OpenStudy (anonymous):

What are the coordinates of the x-intercepts of the graph of y = 2x2 + 6x – 20? (1 point)

OpenStudy (anonymous):

Y = 2x2 + 6x-20. So are you aware of what is an x-intercept?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok. basically x-intercept is a point where the parabola hits x-axis i.e. a spot where y equals 0. So y = 0 Hence 0 = 2x2+6x - 20. divide both sides by 2 - 0 = x2+3x-10 = x2+5x-2x-10 = 0. Solve for x, (x,0) is the x-intercept.

OpenStudy (anonymous):

These are the choices: A. (–5, 0), (2, 0) B. (5, 0), (–2, 0) C. (–4, 0), (10, 0) D. (4, 0), (–10, 0)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so @TrustMeAndWeWillGoFar why dnt u find it out?

OpenStudy (anonymous):

I will, I am working on it.

OpenStudy (anonymous):

So far I have \[20=2x2+6x\]

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

well @TrustMeAndWeWillGoFar i've made that much simpler already. 20 = 2x2+ 6x divide by 2 on both sides - 10 = x2+3x x2+3x-10 = 0 x2+5x-2x-10 = 0 (splitting middle term.) now if u can solve

OpenStudy (anonymous):

hi i need help with Quadrilaterals

OpenStudy (anonymous):

@fun if u can write ur qn separately

OpenStudy (anonymous):

\[x2+5x=2x\]

OpenStudy (anonymous):

What about that?

OpenStudy (anonymous):

Then subtract 5x.

OpenStudy (anonymous):

\[2x=-3x\]

OpenStudy (anonymous):

x2+5x-2x-10 = 0 u can factorise it. x(x+5) -2(x+5) = 0 (x-2)(x+5) = 0 (taking x+5 common) so either x-2 = or x+5 = 0 x = 2, or x = -5

OpenStudy (anonymous):

Ok so it would be \[(2,0) and (-5,0)\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Cool thanks. I'm going to close the question nowe.

OpenStudy (anonymous):

now*